OpenStudy (anonymous):

Suppose this curve y=x^4+ax^3+bx^2+cx+d has a tangent line when x=0 with equation y=2x+1 and a tangent line when x=1 with equation y=6x−3. Find the values of a, b, c, and d.

OpenStudy (anonymous):

First differentiate.

OpenStudy (anonymous):

\(f'(0)=2\) and \(f'(1)=6\)

OpenStudy (anonymous):

so c=0

OpenStudy (anonymous):

No, \(c=2\).

OpenStudy (anonymous):

if x=0 then everything cancels out except for c 4x^3+3ax^2+2bx+c

OpenStudy (dumbcow):

for tangent line: y = mx+ b \[m = f'(x)\]

OpenStudy (anonymous):

Yeah and \(f'(0)=2\) so you end up with \(c=2\).

OpenStudy (anonymous):

ok and a=(-2/3)b

OpenStudy (anonymous):

Okay. \(f(0)=1\)

OpenStudy (anonymous):

and \(f(1)=6(1)-3=3\)

OpenStudy (anonymous):

Because we know tangent lines have to intersect at these points to be tangent.

OpenStudy (anonymous):

ok so

OpenStudy (dumbcow):

set f(1) = 3 and solve for "b" given c=2, d=1, a = (-2/3)b

OpenStudy (anonymous):

b=-3?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

so would i do the same thing to find a

OpenStudy (anonymous):

i got it. a=2

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

thanks @dumbcow and @wio